# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v3

@@ 2.3 Binary Sample
The color-magnitude (CM) diagram for stars is another useful diagnostic for suspicious unresolved close binaries because colors and luminosities of such systems deviate from the the main-sequence locus. Figure 11 shows the CM diagram for the stars in the statistical sample using the Gaia BP--RP color and the $G$-band absolute magnitude $M_{G}$. Most stars fall on a well-defined main sequence. For 22 binaries out of 652, one or both components lie $>2\sigma$ above the scatters of $M_{G}$ at BP--RP. These are potentially suspicious systems as they may be abnormal beyond statistical fluctuations and will not included in inferring gravity.
Figure 12: The left panel shows the profiles of $M_{\rm{tot}}$ as functions of $s/r_{\rm{M}}$, where $s$ is the projected separation and $r_{\rm{M}}$ is the MOND radius (Equation (31)), for the base sample and the statistical sample listed in Table 1. While both profiles are flat, the median is much higher in the statistical sample. This indicates that stars with more precisely measured RVs are likely to be brighter. The middle panel shows only the profile of $v_{p}$ for the base sample as the measured $v_{r}$ values are too imprecise for general wide binaries (see Figure 9). It can be seen that $v_{p}$ is boosted with respect to the Keplerian extrapolation in the last three bins consistent with the findings of Chae (2024a, c) and Hernandez et al. (2024a). The right panel shows the profiles of the three velocity components for the binaries in the statistical sample with the limit ruwe $<1.3$. In the last bin of the MOND regime ($s/r_{\rm{M}}>0.5$), boosts of the three velocity components can be seen. But, notice that the medians of $|v_{r}|$ in the middle and last bins are markedly larger than $|v_{x^{\prime}}|$ and $|v_{y^{\prime}}|$ that are statistically indistinguishable each other. The additional boost can be understood as a consequence of the relatively large measurement errors of RVs.
For the binary stars of the selected samples stellar masses are estimated using the standard $M_{G}$-mass relation obtained by Chae (2023a) (the first choice in Table 1 of that work) where $M_{G}$ is the $G$-band absolute magnitude corrected for dust extinction as described in Chae (2023a). The left panel of Figure 13 shows the distribution of the Chae (2023a) masses of 1304 stars in the statistical sample. For these masses that are mainly $\gtrsim 0.5M_{\odot}$, the uncertainties of 5% may well encompass alternative masses as shown in Figure 7 of Chae (2023a). For alternative masses, it is relevant to consider those provided by the Gaia DR3 Final Luminosity Age Mass Estimator (FLAME). FLAME masses are independent and individual results for selective stars from stellar evolution modeling based on a host of Gaia photometric, spectroscopic, and astrometric data. FLAME masses are available for only 859 out of the 1304 stars. The right panel of Figure 13 shows the ratio of the FLAME mass to the Chae (2023a) mass for the subset of stars. The mean and standard deviation of the ratios are $0.97$ and $0.05$ indicating that two masses are indeed consistent with each other up to about 5% scatter. Because FLAME masses are not available for all stars, masses in this work refer to the Chae (2023a) values unless stated otherwise.
@@
where $a_{0}=1.2\times 10^{-10}$ m s-2 is estimated by observations (e.g. McGaugh et al. 2016). The left panel of Figure 12 shows that $M_{\rm{tot}}$ does not vary statistically with $s/r_{\rm{M}}$. This means that Newtonian gravity would predict a statistically Keplerian velocity profile with $s/r_{\rm{M}}$ as long as eccentricity does not appreciably vary with $s/r_{\rm{M}}$ in a statistical sense. The middle panel shows the profile of $v_{p}$ for the base sample. Note that all $v_{p}$ values from the base sample have very small errors and thus all of them are used to obtain the profile. It agrees well with the Keplerian profile of $v_{p}\propto s^{-1/2}$ for the first four bins but indicates some boost with respect to the Keplerian prediction for the last three bins. This is in line with the finding of Chae (2024a, c).
@@ 2.4 Mock binary samples
In Section 2.2 each set of mock observations (i.e. a mock sample) are generated for one mock binary with specific parameters. For comparison with the samples of snapshot observed wide binaries that have different parameters, here I describe a procedure to generate samples of mock snapshot observed wide binaries satisfying Newtonian or pseudo-Newtonian gravity with a generalized gravitational constant $G=\gamma_{g}G_{\rm{N}}$.
As in Chae (2023a, 2024a, 2024c), mock binaries are generated using the observed binaries by replacing the observed positions, PMs, and RVs with simulated values as described below. Because all three components of the relative velocity between the pair need to be generated in the observer’s frame, I consider an arbitrary transformation from the orbital plane to the observer’s frame with Euler angles as described in Appendix A. Since what matter are the relative positions and motions between the pair, the positions and velocities of the component $A$ will be fixed at the observed values.
@@
    The sky-projected separation $s$ for each binary system is the same as the Gaia measured value. The distance of the system from the Sun is taken to be the error-weighted mean ($d_{M}$) of the Gaia measured values for the two components. The total mass $M_{\rm{tot}}$ is the same as the estimated value as described in Section 2.3. The eccentricity ($e$) of the orbit is taken randomly from the Hwang et al. (2022) maximum likelihood range for the binary, as described in Section 2.4 of Chae (2024c).
@@
    In principle, three Euler angles ($\alpha,\beta,\gamma$) defined in Appendix A need to be assigned. The first and third angles ($\alpha$ and $\gamma$) represent rotations about the initial and final “$z$” axes. However, when $\phi_{0}$ is randomly taken as is the case here, $\alpha$ can be fixed at zero. Let $\gamma=\theta$ with $\theta$ taken randomly from $(-\pi,\pi)$. The second angle $\beta$ represents a rotation about the intermediate “$x$” axis. This is an inclination as it has been referred to so. Let $\beta=i$ with $i=\cos^{-1}(u)$ where $u$ is taken randomly from $(0,1)$ so that the probability density of $i$ is $\sin i$. The two angles $i$ and $\theta$ (with a random assignment of $\phi_{0}$) specify a general transformation matrix $\mathbf{\Lambda}=\mathbf{\Lambda}(i,\theta)$ (Equation (A2)) between the “$xyz$” frame (with the orbit lying on its $xy$ plane) and the observer’s “$x^{\prime}y^{\prime}z^{\prime}$” frame.
@@
    The mock relative R.A. and decl. of component $B$ with respect to $A$ are given by $\Delta\alpha=-x^{\prime}/d_{M}$ and $\Delta\delta=y^{\prime}/d_{M}$ in units of radian. These are converted to units of degree to mimic the Gaia data. Although $z^{\prime}$ represents the difference in the distances of the two components, it has no use in the present simulations because the Gaia measurement errors are typically much larger than $|z^{\prime}|$.
@@
    The velocity component $v_{z^{\prime}}$ (i.e. $v_{r}$) gives the relative RV of component $B$ with respect to $A$ by reversing the left- and right-hand sides of Equation (7). These assigned values of RVA and RVB have no errors. However, unlike the Gaia reported PMs giving accurate values of $v_{x^{\prime}}$ and $v_{y^{\prime}}$, the Gaia reported errors of RV values are non-negligible. In other words, their errors mean $\sigma_{v_{r}}\gtrsim 170$ m s-1 (Figure 9) for $|v_{r}|=|\rm{RV}_{B}-\rm{RV}_{A}|$ which can be comparable to, or even larger than, $|v_{z^{\prime}}|$ for wide binaries with relatively large separation ($s\gtrsim 5$ kau). Such large errors can bias $|v_{r}|$ statistically. Thus, the model predicted RVA and RVB values are given Gaussian scatters with the Gaia measurement errors (or other errors as needed by the desired test).
@@
A sample of mock Newtonian binaries generated for the base sample shown in Figure 12 can be found in Figure 14. The left panel shows the profile of $v_{p}(=\sqrt{v_{x^{\prime}}^{2}+v_{y^{\prime}}^{2}})$, which agrees well with the Keplerian profile of $v_{p}\propto s^{-1/2}$. The middle and right panels show the profiles of the three velocity components $|v_{x^{\prime}}|$, $|v_{y^{\prime}}|$, and $|v_{z^{\prime}}|$ in mock Newtonian samples generated without and with RV scatters for the statistical sample. The right panel clearly reveals the effects of RV scatters. While $|v_{x^{\prime}}|$ and $|v_{y^{\prime}}|$ follow the Keplerian profile, $|v_{z^{\prime}}|$ deviates increasingly at larger $s/r_{\rm{M}}$ due to the relatively larger effects of RV scatters for smaller $|v_{r}|$. Nevertheless, note that a comparison between the right panel of Figure 12 and that of Figure 14 indicates a boost in the ‘MOND regime’ bin of the Gaia sample.
@@ 3.1 Overview and inspection of the individual Bayesian results
In all cases the individual PDFs of $\Gamma$ are much broader than the difference $\approx 0.07$ between Newton and MOND, and thus any specific result strongly inconsistent (e.g. $>3\sigma$) with the range $0\leq\Gamma\leq 0.07$ set by Newton and MOND may signal kinematic contaminants of hidden low-mass objects (e.g. faint stars or very massive Jovian planets) close to either or both of the stars, or some unknown issues with the RV data999 See Katz et al. (2023) for a general description of the Gaia DR3 RV data.. Although the statistical sample was selected through a rigorous screening process, rare exceptions may be present. As a specific example, consider an unobserved companion star with $0.1M_{\odot}$ orbiting a host star with $1.0M_{\odot}$ at a separation of 10 au. For the sake of simplicity, assume a circular orbit. Then, the relative velocity is $v\approx 9900$ m s-1 according to Equation (23) and the velocity perturbation experienced by the host star is $\approx 9900\times 0.1/1.1=900$ m s-1, which is significant. When the measured relative velocity between the two main stars is boosted by such a perturbation, the Bayesian modeling may return a PDF of $\Gamma$ that excludes the range $0\leq\Gamma\leq 0.07$ with a high significance. This is confirmed with a mock sample including hierarchical systems.
@@
Note. (1) The quoted values of $e$ and $\Gamma$ are from the posterior PDFs as examples are given in Figure 16. If these were pure wide binaries without hidden kinematic contaminants and the RV values and uncertainties were reliable, one of these systems alone would rule out Newtonian gravity. (2) Most stars have normal ruwe values; ${\tt ruwe}<1.3$ are satisfied by all except for three stars, which have $1.3\lesssim{\tt ruwe}\lesssim 1.5$ that are still not so large. This indicates that kinematically abnormal systems based on Gaia data not necessarily have large ruwe values.
Any kinematic perturbers may cause outliers only in the positive direction of $\Gamma$. This is indeed the case for the Bayesian results of the statistical sample. To be concrete, I define the lower ($\sigma_{\rm{low}}$) and upper ($\sigma_{\rm{up}}$) uncertainties of $\Gamma$ with the 16th and 84th percentiles in its PDF. Then, $\Gamma_{\rm{med}}+3\sigma_{\rm{up}}>0$ is satisfied in all cases, but $\Gamma_{\rm{med}}-3\sigma_{\rm{down}}<0.02$ (here $0.02$ is used to allow for a median effect of RV uncertainties in a transition regime between the Newtonian and the MOND regimes) is not satisfied by eight wide binaries. Two examples of these outliers can be found in Figure 16. Clearly, any of these results alone would not only rule out the range $0\leq\Gamma\leq 0.07$, but would also be inconsistent with the PDFs of many other wide binaries. The properties of the eight wide binaries with suspected kinematic contaminants or abnormal RV data are summarized in Table 2. These systems will not be included in the inferring of consolidated gravity in the following subsections. In doing so, it is guaranteed that all individual binaries are consistent with Newtonian gravity within $\approx 3\sigma$ and the consolidated gravity is not dominated by a few radically different ones, but determined by the collective trend of the majority. See, however, Section 4.1 for discussions on alternative interpretations.
@@
Properties of the inferred orbital and geometric parameters are of interest not only to check whether they are reasonable but also in its own right. Eccentricities of wide binaries are of particular interest. For all wide binaries in the El-Badry et al. (2021) catalog from which the statistical sample was drawn, Hwang et al. (2022) have inferred eccentricities with an independent Bayesian method using only the sky-projected relative displacements and relative velocities (without RVs). Figure 17 compares the eccentricities from this work with those from Hwang et al. (2022). The eccentricities with the thermal prior agree excellently with the Hwang et al. (2022) values as those with the H2022 prior do. However, those with the flat prior show some discrepancy. This is not surprising given that the statistical properties of observationally inferred eccentricities for wide binaries (see Figure 24 of Chae (2023a)) are much more consistent with the thermal or H2022 distribution than the flat distribution. Moreover, as will be shown in Section 3.2 below, a test with the Newtonian control sample indicates that the flat prior can cause some bias in the Bayesian inference of gravity when RVs are not sufficiently precise. The right column of Figure 17 compares the widths of the PDFs between this work and Hwang et al. (2022). Because this work uses an additional constraint from RVs, it is expected that the widths from this work will be narrower. Indeed, this is the case. In particular, the results with the thermal or H2022 prior exhibit narrower widths more clearly.
@@
For elliptical orbits, random snapshot observations will not have a uniform occurrence rate of $\phi-\phi_{0}$ where $\phi$ is the observed phase $\phi=\tan^{-1}((\Delta y^{\prime}/\Delta x^{\prime})/\cos i_{\rm{obs}})$ with $i_{\rm{obs}}=\tan^{-1}(-v_{r}/v_{y^{\prime}})$ and $\phi_{0}$ is the longitude of the periastron. Thus, it is interesting to check the posterior distribution of $\phi-\phi_{0}$ for the statistical sample where $\phi_{0}$ is drawn from the PDF of each wide binary. Figure 18 shows the distribution of $\phi-\phi_{0}$ for the 652 wide binaries of the statistical sample. The histogram represents a median of many distributions where each distribution represents values drawn once from all wide binaries. The error bars represent the scatters from one draw to another. This posterior distribution agrees remarkably well with the prior distribution given by Equation (19) with a reasonable representative value of $e=0.76_{-0.06}^{+0.05}$, indicating that snapshot observations are truly random in normalized time $t/P$ (where $t$ is the time from the passage of the periastron and $P$ is the orbital period).
@@
The posterior distribution of $\log_{10}f_{M}$ always agrees well with the prior input distribution because no observational constraints have any dependence on it. Allowing for a finite width of its PDF rather than fixing $\log_{10}f_{M}=0$ serves merely as a way to take into account uncertainties of mass in the inference of $\Gamma$.
@@
Since the goal of this work is to test gravity in low acceleration regimes ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2), binaries are split by their internal acceleration between the pair. Binaries in a relatively strong acceleration regime ($g_{\rm{N}}\gtrsim 10^{-8}$ m s-2) are considered in Section 3.2 while binaries in low acceleration regimes are considered in Section 3.3.

### v4

@@ 2.3 Binary Sample
The color-magnitude (CM) diagram for stars is another useful diagnostic for suspicious unresolved close binaries because colors and luminosities of such systems deviate from the the main-sequence locus. Figure 11 shows the CM diagram for the stars in the statistical sample using the Gaia BP--RP color and the $G$-band absolute magnitude $M_{G}$. Most stars fall on a well-defined main sequence. For 22 binaries out of 652, one or both components lie $>2\sigma$ above the scatters of $M_{G}$ at BP--RP. These are potentially suspicious systems as they may be abnormal beyond statistical fluctuations and will not be included in inferring gravity.
For the binary stars of the selected samples stellar masses are estimated using the standard $M_{G}$-mass relation obtained by Chae (2023a) (the first choice in Table 1 of that work) where $M_{G}$ is the $G$-band absolute magnitude corrected for dust extinction as described in Chae (2023a). The left panel of Figure 12 shows the distribution of the Chae (2023a) masses of 1304 stars in the statistical sample. For those masses that are mainly $\gtrsim 0.5M_{\odot}$, the uncertainties of 5% may well encompass alternative masses as shown in Figure 7 of Chae (2023a). For alternative masses, it is relevant to consider those provided by the Gaia DR3 Final Luminosity Age Mass Estimator (FLAME). FLAME masses are independent and individual results for selective stars from stellar evolution modeling based on a host of Gaia photometric, spectroscopic, and astrometric data. FLAME masses are available for only 859 out of the 1304 stars. The right panel of Figure 12 shows the ratio of the FLAME mass to the Chae (2023a) mass for the subset of stars. The mean and standard deviation of the ratios are $0.97$ and $0.05$ indicating that the two masses are indeed consistent with each other up to about 5% scatter. Because FLAME masses are not available for all stars, the masses in this work refer to the Chae (2023a) values unless stated otherwise.
@@
Figure 13: The left panel shows the profiles of $M_{\rm{tot}}$ as a function of $s/r_{\rm{M}}$, where $s$ is the projected separation and $r_{\rm{M}}$ is the MOND radius (Equation (31)), for the base sample and the statistical sample listed in Table 1. While both profiles are flat, the median is much higher in the statistical sample. This indicates that stars with more precisely measured RVs are likely to be brighter. The middle panel shows only the profile of $v_{p}$ for the base sample as the measured $v_{r}$ values are too imprecise for general wide binaries (see Figure 9). It can be seen that $v_{p}$ is boosted with respect to the Keplerian extrapolation in the last three bins consistent with the findings of Chae (2024a, c) and Hernandez et al. (2024a). The right panel shows the profiles of the three velocity components for the binaries in the statistical sample with the limit ruwe $<1.3$. In the last bin of the MOND regime ($s/r_{\rm{M}}>0.5$), boosts of the three velocity components can be seen. But, notice that the medians of $|v_{r}|$ in the middle and last bins are markedly larger than $|v_{x^{\prime}}|$ and $|v_{y^{\prime}}|$, which are statistically indistinguishable each other. The additional boost can be understood as a consequence of the relatively large measurement errors of the RVs.
@@
where $a_{0}=1.2\times 10^{-10}$ m s-2 is estimated by observations (e.g. McGaugh et al. 2016). The left panel of Figure 13 shows that $M_{\rm{tot}}$ does not vary statistically with $s/r_{\rm{M}}$. This means that Newtonian gravity would predict a statistically Keplerian velocity profile with $s/r_{\rm{M}}$ as long as the eccentricity does not appreciably vary with $s/r_{\rm{M}}$ in a statistical sense. The middle panel shows the profile of $v_{p}$ for the base sample. Note that all $v_{p}$ values from the base sample have very small errors and thus all of them are used to obtain the profile. It agrees well with the Keplerian profile of $v_{p}\propto s^{-1/2}$ for the first four bins but indicates some boost with respect to the Keplerian prediction for the last three bins. This is in line with the findings of Chae (2024a, c).
@@ 2.4 Mock binary samples
In Section 2.2 each set of mock observations (i.e., a mock sample) are generated for one mock binary with specific parameters. For comparison with the samples of snapshot-observed wide binaries that have different parameters, here I describe a procedure to generate samples of mock snapshot-observed wide binaries satisfying Newtonian or pseudo-Newtonian gravity with a generalized gravitational constant $G=\gamma_{g}G_{\rm{N}}$.
As in Chae (2023a, 2024a, 2024c), mock binaries are generated using the observed binaries by replacing the observed positions, PMs, and RVs with simulated values as described below. Because all three components of the relative velocity between the pair need to be generated in the observer’s frame, I consider an arbitrary transformation from the orbital plane to the observer’s frame with Euler angles as described in Appendix A. Since what matters are the relative positions and motions between the pair, the positions and velocities of the component $A$ will be fixed at the observed values.
@@
    The sky-projected separation $s$ for each binary system is the same as the Gaia-measured value. The distance of the system from the Sun is taken to be the error-weighted mean ($d_{M}$) of the Gaia-measured values for the two components. The total mass $M_{\rm{tot}}$ is the same as the estimated value as described in Section 2.3. The eccentricity ($e$) of the orbit is taken randomly from the Hwang et al. (2022) maximum likelihood range for the binary, as described in Section 2.4 of Chae (2024c).
@@
    In principle, the three Euler angles ($\alpha,\beta,\gamma$) defined in Appendix A need to be assigned. The first and third angles ($\alpha$ and $\gamma$) represent rotations about the initial and final “$z$-axes”. However, when $\phi_{0}$ is randomly taken as is the case here, $\alpha$ can be fixed at zero. Let $\gamma=\theta$ with $\theta$ taken randomly from $(-\pi,\pi)$. The second angle $\beta$ represents a rotation about the intermediate “$x$-axis”. This is an inclination as it has been referred to. Let $\beta=i$ with $i=\cos^{-1}(u)$ where $u$ is taken randomly from $(0,1)$ so that the probability density of $i$ is $\sin i$. The two angles $i$ and $\theta$ (with a random assignment of $\phi_{0}$) specify a general transformation matrix $\mathbf{\Lambda}=\mathbf{\Lambda}(i,\theta)$ (Equation (A2)) between the “$xyz$” frame (with the orbit lying on its $xy$ plane) and the observer’s “$x^{\prime}y^{\prime}z^{\prime}$” frame.
@@
    The mock relative R.A. and decl. of component $B$ with respect to $A$ are given by $\Delta\alpha=-x^{\prime}/d_{M}$ and $\Delta\delta=y^{\prime}/d_{M}$ in units of radians. These are converted to units of degrees to mimic the Gaia data. Although $z^{\prime}$ represents the difference in the distances of the two components, it has no use in the present simulations because the Gaia measurement errors are typically much larger than $|z^{\prime}|$.
@@
    The velocity component $v_{z^{\prime}}$ (i.e., $v_{r}$) gives the relative RV of component $B$ with respect to $A$ by reversing the left- and right-hand sides of Equation (7). These assigned values of RVA and RVB have no errors. However, unlike the Gaia-reported PMs giving accurate values of $v_{x^{\prime}}$ and $v_{y^{\prime}}$, the Gaia-reported errors of RV values are nonnegligible. In other words, their errors mean $\sigma_{v_{r}}\gtrsim 170$ m s-1 (Figure 9) for $|v_{r}|=|\rm{RV}_{B}-\rm{RV}_{A}|$ which can be comparable to, or even larger than, $|v_{z^{\prime}}|$ for wide binaries with relatively large separation ($s\gtrsim 5$ kau). Such large errors can bias $|v_{r}|$ statistically. Thus, the model-predicted RVA and RVB values are given Gaussian scatter with the Gaia measurement errors (or other errors as needed by the desired test).
@@
A sample of mock Newtonian binaries generated for the base sample shown in Figure 13 can be found in Figure 14. The left panel shows the profile of $v_{p}(=\sqrt{v_{x^{\prime}}^{2}+v_{y^{\prime}}^{2}})$, which agrees well with the Keplerian profile of $v_{p}\propto s^{-1/2}$. The middle and right panels show the profiles of the three velocity components $|v_{x^{\prime}}|$, $|v_{y^{\prime}}|$, and $|v_{z^{\prime}}|$ in mock Newtonian samples generated without and with RV scatter for the statistical sample. The right panel clearly reveals the effects of RV scatter. While $|v_{x^{\prime}}|$ and $|v_{y^{\prime}}|$ follow a Keplerian profile, $|v_{z^{\prime}}|$ deviates increasingly at larger $s/r_{\rm{M}}$ due to the relatively larger effects of RV scatter for smaller $|v_{r}|$. Nevertheless, note that a comparison between the right panel of Figure 13 and that of Figure 14 indicates a boost in the “MOND regime” bin of the Gaia sample.
@@ 3.1 Overview and inspection of the individual Bayesian results
In all cases the individual PDFs of $\Gamma$ are much broader than the difference $\approx 0.07$ between Newton and MOND, and thus any specific result strongly inconsistent (e.g., $>3\sigma$) with the range $0\leq\Gamma\leq 0.07$ set by Newton and MOND may signal kinematic contaminants of hidden low-mass objects (e.g., faint stars or very massive Jovian planets) close to either or both of the stars, or some unknown issues with the RV data101010 See Katz et al. (2023) for a general description of the Gaia DR3 RV data.. Although the statistical sample was selected through a rigorous screening process, rare exceptions may be present. As a specific example, consider an unobserved companion star with $0.1M_{\odot}$ orbiting a host star with $1.0M_{\odot}$ at a separation of 10 au. For the sake of simplicity, assume a circular orbit. Then, the relative velocity is $v\approx 9900$ m s-1 according to Equation (23) and the velocity perturbation experienced by the host star is $\approx 9900\times 0.1/1.1=900$ m s-1, which is significant. When the measured relative velocity between the two main stars is boosted by such a perturbation, the Bayesian modeling may return a PDF of $\Gamma$ that excludes the range $0\leq\Gamma\leq 0.07$ with a high significance. This is confirmed with a mock sample including hierarchical systems.
@@
Note. (1) The quoted values of $e$ and $\Gamma$ are from the posterior PDFs: examples are given in Figure 16. If these were pure wide binaries without hidden kinematic contaminants and the RV values and uncertainties were reliable, one of these systems alone would rule out Newtonian gravity. (2) Most stars have normal ruwe values; ${\tt ruwe}<1.3$ is satisfied by all except for three stars, which have $1.3\lesssim{\tt ruwe}\lesssim 1.5$ that are still not so large. This indicates that kinematically abnormal systems based on Gaia data not necessarily have large ruwe values.
Any kinematic perturbers may cause outliers only in the positive direction of $\Gamma$. This is indeed the case for the Bayesian results of the statistical sample. To be concrete, I define the lower ($\sigma_{\rm{low}}$) and upper ($\sigma_{\rm{up}}$) uncertainties of $\Gamma$ with the 16th and 84th percentiles in its PDF. Then, $\Gamma_{\rm{med}}+3\sigma_{\rm{up}}>0$ is satisfied in all cases, but $\Gamma_{\rm{med}}-3\sigma_{\rm{down}}<0.02$ (here $0.02$ is used to allow for a median effect of RV uncertainties in a transition regime between the Newtonian and the MOND regimes) is not satisfied by eight wide binaries. Two examples of these outliers can be found in Figure 16. Clearly, any of these results alone would not only rule out the range $0\leq\Gamma\leq 0.07$, but would also be inconsistent with the PDFs of many other wide binaries. The properties of the eight wide binaries with suspected kinematic contaminants or abnormal RV data are summarized in Table 2. These systems will not be included in the inference of consolidated gravity in the following subsections. In doing so, it is guaranteed that all individual binaries are consistent with Newtonian gravity within $\approx 3\sigma$ and the consolidated gravity is not dominated by a few radically different ones, but determined by the collective trend of the majority. See, however, Section 4.1 for discussions on alternative interpretations.
@@
Properties of the inferred orbital and geometric parameters are of interest not only to check whether they are reasonable but also in their own right. Eccentricities of wide binaries are of particular interest. For all wide binaries in the El-Badry et al. (2021) catalog from which the statistical sample was drawn, Hwang et al. (2022) have inferred eccentricities with an independent Bayesian method using only the sky-projected relative displacements and relative velocities (without RVs). Figure 17 compares the eccentricities from this work with those from Hwang et al. (2022). The eccentricities with the thermal prior agree excellently with the Hwang et al. (2022) values, as do those with the H2022 prior. However, those with the flat prior show some discrepancy. This is not surprising given that the statistical properties of observationally inferred eccentricities for wide binaries (see Figure 24 of Chae (2023a)) are much more consistent with the thermal or H2022 distribution than the flat distribution. Moreover, as will be shown in Section 3.2 below, a test with the Newtonian control sample indicates that the flat prior can cause some bias in the Bayesian inference of gravity when the RVs are not sufficiently precise. The right column of Figure 17 compares the widths of the PDFs between this work and Hwang et al. (2022). Because this work uses an additional constraint from RVs, it is expected that the widths from this work will be narrower. Indeed, this is the case. In particular, the results with the thermal or H2022 prior exhibit narrower widths more clearly.
@@
For elliptical orbits, random snapshot observations will not have a uniform occurrence rate of $\phi-\phi_{0}$ where $\phi$ is the observed phase $\phi=\tan^{-1}((\Delta y^{\prime}/\Delta x^{\prime})/\cos i_{\rm{obs}})$ with $i_{\rm{obs}}=\tan^{-1}(-v_{r}/v_{y^{\prime}})$ and $\phi_{0}$ is the longitude of the periastron. Thus, it is interesting to check the posterior distribution of $\phi-\phi_{0}$ for the statistical sample where $\phi_{0}$ is drawn from the PDF of each wide binary. Figure 18 shows the distribution of $\phi-\phi_{0}$ for the 652 wide binaries of the statistical sample. The histogram represents the median of many distributions where each distribution represents values drawn once from all wide binaries. The error bars represent the scatter from one draw to another. This posterior distribution agrees remarkably well with the prior distribution given by Equation (19) with a reasonable representative value of $e=0.76_{-0.06}^{+0.05}$, indicating that snapshot observations are truly random in normalized time $t/P$ (where $t$ is the time from the passage of the periastron and $P$ is the orbital period).
@@
The posterior distribution of $\log_{10}f_{M}$ always agrees well with the prior input distribution because no observational constraints have any dependence on it. Allowing for a finite width of its PDF rather than fixing $\log_{10}f_{M}=0$ serves merely as a way to take into account the uncertainties of mass in the inference of $\Gamma$.
@@
Since the goal of this work is to test gravity in low-acceleration regimes ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2), binaries are split by their internal acceleration between the pair. Binaries in a relatively strong-acceleration regime ($g_{\rm{N}}\gtrsim 10^{-8}$ m s-2) are considered in Section 3.2 while binaries in low-acceleration regimes are considered in Section 3.3.